OpenStudy (anonymous):

URGENT HELP!! WILL FAN AND GIVE MEDAL!! Solve the following equation:

OpenStudy (anonymous):

\[16^x+1024=5 \times 4^x + 8\]

jhonyy9 (jhonyy9):

little help 16^x =4^2x

jhonyy9 (jhonyy9):

1024=4^5